OpenStudy (anonymous):

The range of the function y = cos x + 1 is -1 ≤ y ≤ 1 True or false?

OpenStudy (anonymous):

What kind of values do you think cos(x) can produce?

OpenStudy (anonymous):

idk?

OpenStudy (anonymous):

I'd suggest having a look at the theory behind trigonometric functions. However, to help you with this issue, \[-1 \leq \cos(x) \leq 1\]